# Masked Image Modeling via Dynamic Token Morphing

### v1

@@ 4.2 Dynamic Token Morphing
At this point, the token morphing function for the -th level contextual representations can be determined by . Note that each -th level token morphing function is dynamically changed based on randomly sampled and . Eventually, we can obtain the -th level token morphing matrix , a contextual relation among tokens, from the target token representation as follows:
where indicates the aggregation of the -th token to the -th token . In addition, every token is assigned to a specific cluster, even in cases where it forms a single token cluster, and each token retains its exclusive association with a single cluster. Thus, morphing matrix should satisfies the following two conditions, and . Our DTM is universal to any token aggregation method; we simply employ bipartite matching to efficiently search for contextually similar tokens.
@@
Based on the -th level token morphing matrix obtained from dynamic token morphing, the representations for online and target morphed tokens can be derived by:
where and denote for the representations of online and target morphed tokens, respectively.
| Method | ViT-S | ViT-B | ViT-L |
| MIM (Baseline) | 82.7 | 84.3 | 86.4 |
| DTM (Ours) | 83.1 | 85.4 | 86.7 |
| Gain | (+0.4) | (+1.1) | (+0.3) |
| Method | Pre-training epochs | Supervision | ViT-S | ViT-B | ViT-L | Seg. | |||
@@
| SplistMask [14] | arXiv 2021 | 300 | 300 | - | Pixel + Feature | 81.5 | 83.6 | - | 46.8 |
@@
| CAE [10] | arXiv 2022 | 300 | 1600 | 1600 | Pixel + Feature | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE [18] | arXiv 2022 | - | 1600 | 1600 | Pixel + Feature | - | 84.4 | 85.3 | 50.1 |
@@
| BootMAE [12] | ECCV 2022 | - | 800 | 800 | Pixel + Feature | - | 84.2 | 85.9 | 49.1 |
| CAN [25] | arXiv 2023 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
@@
4.3 Training Objective
We formulate the objective function by accumulating DTM loss, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss can be formulated as follows:
where is the i-th distance function, is a number of aggregated tokens for the -th online and target morphed tokens. Note that our DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in Eq. 4 for all the DTM losses. Therefore, the final objective function is the summation of all DTM losses, which is defined as:
@@ 5 Experiment
Pre-training setup.
We use the training dataset of ImageNet-1K for pre-training. We pre-train our framework and other SSL frameworks through DTM on ImageNet-1K. We pre-train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base model with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4. Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. We also pre-train various SSL frameworks through our DTM. Detailed pre-training setups can be found in Appendix. All models are pre-trained using 8 V100-32GB GPUs.
| Method | iNat 2018 | iNat 2019 | iNat 2021-mini |
| Baseline | 75.0 (74.6±0.6) | 81.1 (79.8±1.0) | 75.8 (75.2±0.6) |
| DTM (Ours) | 78.5 (77.4±0.6) | 81.9 (81.2±0.6) | 78.4 (77.5±0.6) |
Fine-tuning setup.
We fine-tune our pre-trained models on ImageNet-1K [34] by default. Pre-trained ViT-S, ViT-B, and ViT-L are fine-tuned for 300, 100, and 50 epochs, respectively. The pre-trained models are optimized by AdamW with a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. Additional fine-tuning details are listed in Appendix.
5.1 Effectiveness of Our Method
Applicability.
We apply DTM on BEiT v2, MAE with the CLIP teacher [29], and BYOL [15] to verify its applicability to various SSL frameworks. As shown in Table 3, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability on various frameworks beyond our baseline.
Scalability.
We further study the scalability of DTM. We compare the fine-tuning accuracies of our method to its baseline, which is pre-trained without DTM. We pre-train ViT-S/16, ViT-B/16, and ViT-L/16 with a resolution of for each method and evaluate the performance gains in fine-tuning accuracies. As shown in Table 5, DTM improved ViT-S/16, ViT-B/16, and ViT-L/16 by 0.4p, 1.1p, and 0.3p, respectively, verifying its general efficacy across different scales.
| Method | Dynamic | FT. () | |
@@ 5.1 Effectiveness of Our Method
5.2 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K [34] datasets. The comparisons include supervised learning and SSL methods with various supervision signals. In the case when the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for fair comparison. Table 5 reports the fine-tuning accuracies of ViT-S/B/L backbones. We observe that our MIM pre-trained by DTM achieves 83.2, 85.4, 86.7 top-1 accuracy with ViT-S/16, ViT-B/16, and ViT-L/16, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP, DeepMIM, and BEiT v2 by 1.0p, 0.6 p, and 0.4p on ViT-B/16, respectively. Moreover, our method definitely outperforms other methods that leverage supervisions other than CLIP representations, which demonstrates the superiority of our method among self-supervised learning methods.
5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performances on ADE20K [50] to verify the transferability of our pre-trained model to dense prediction tasks. We follow the training and evaluation protocol [17]; the models are fine-tuned for 160K iterations using UperNet [46] with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 5 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method is effective in enhancing discriminability for dense prediction tasks.
5.4 Transfer learning
iNaturalist datasets.
Fine-Grained Visual Classification (FGVC) datasets.
@@ 5.5 Analysis
Ablation study on the importance of dynamics.
Table 8 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ a ViT-B/16 model with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K [34]. For token morping without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 8, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, which demonstrates the importance of dynamic mechanism in DTM.
Faster convergence.
We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 4(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 4(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss compared to the baseline model, verifying better convergence. Finally, Fig. 4(c) verifies consistent effect of DTM across various pre-training epochs.
@@ 6 Conclusion
We have introduced a novel masked image modeling method with morphed tokens to address the negative impacts of spatially inconsistent target representations during pre-training. We first analyze the existence and impacts of spatial inconsistency in target representations. Specifically, we qualitatively observed spatial inconsistency among patches from pre-trained models despite proximity and contextual similarity. Then, we investigate the impacts on downstream tasks, including zero-shot image classification and linear classification. We further validate that token morphing enhances the pre-training capability of the target patches, while arbitrary aggregation, like downsampling, tends to disrupt them. The results demonstrate the necessity of token morphing to alleviate spatial inconsistency. Based on the observation, we have proposed Dynamic Token Morphing, dynamically aggregating contextually associated tokens with a random number of matching tokens. Subsequently, our method aligns the representations of morphed tokens with the representations of their corresponding morphed targets. The extensive experiments verify its general applicability, scalability, and superiority. We further validate its transferability on the iNaturalist datasets and Fine-grained Visual Classification datasets. Subsequent analysis reveals superiority over the baseline.
Limitation
@@
Despite the potential of our method, we have verified its applicability only up to ViT-L/16. Resource limitations prevented us from performing more experiments on larger-scale models such as ViT-G. We will explore this in our future work.
Supplementary Material
@@
This supplementary material includes additional experimental analyses of our proposed method. We first present additional examples of spatial inconsistency among patches from pre-trained models. Then, we verify the further applicability of Dynamic Token Morphing (DTM) on SSL framework beyond Masked Image Modeling methods and target models beyond CLIP [29]. Moreover, we conduct an ablation study on loss functions. Finally, we provide implementation details for both pre-training and fine-tuning on ImageNet-1K [34], as well as fine-tuning on ADE20K [50].
Appendix A More examples of spatial inconsistency
@@
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize prediction results for each patch in the bounding boxes through patch-wise zero-shot classification with and without token morphing using CLIP [29]. Specifically, we color the patches in green or red to denote that the predicted classes are related or unrelated to the ground-truth class, respectively. Moreover, we gradually vary the lightness of the colors to exhibit the degree of similarity and dissimilarity. As shown in Fig. 6, patch representations without token morphing reveal spatially inconsistent patch-wise predictions compared to the prediction results with token morphing, which reveals the spatially inconsistent representations among patches. In addition, predictions with token morphing exhibit significantly more accurate patches compared to predictions without token morphing.
Appendix B Applicability on SLIP
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
| Case | FT. () |
| 83.2 | |
| 83.3 | |
| Smoothed | 83.2 |
| Cosine distance (Cos) | 84.2 |
| Case | Throughput (image/s) | FT. () |
| Baseline | 1460.0 | 84.3 |
| Layer-wise K-Means clustering | 1265.4 | 85.1 |
| DTM (K-Means clustering) | 489.1 | 85.4 |
| DTM (Bipartite matching) | 1446.0 | 85.4 |
Appendix C Ablation study on loss functions
We compare various loss functions for our method in Table 10. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [34] While all loss functions work well, the cosine distance works best.
| Hyperparameters | Small Size | Base Size | Large Size |

### v2

@@ 4.2 Overall Pipeline of Dynamic Token Morphing
Eventually, we obtain the token morphing matrix , a contextual relation among tokens, from the target token representations as follows:
where indicates that the token will be aggregated to the token . In addition, every token should be assigned to a specific cluster, even in cases where it forms a single token cluster, and each token should retain its exclusive association with a single cluster (i.e., and ). For the repetition , token pairs are successively matched.
@@
Based on the token morphing matrix obtained by , the representations for online and target morphed tokens are derived by:
where and denote for the representations of online and target morphed tokens, respectively. Note that the morphed tokens are representative tokens for each token group, with their representations being smoothed specific to their respective groups.
4.2.3 Aligning morphed tokens.
We formulate the objective function by accumulating DTM losses, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss with sampled and is formulated as follows:
where is a distance function and is a number of tokens aggregated for the online and target morphed tokens. Here, we utilize to consider all tokens aggregated for the morphed tokens. Note that our DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in Eq. 4 for all the DTM losses. To further enhance the dynamism of our method, we concurrently apply multiple DTMs. Therefore, the final objective function is the summation of all DTM losses, which is defined as:
where denotes the total number of simultaneously employed DTMs.
| Method | Pre-training epochs | Supervision | ViT-S | ViT-B | ViT-L | Seg | |||
@@
| SplistMask [16] | arXiv 2021 | 300 | 300 | - | Pixel+Feat | 81.5 | 83.6 | - | 46.8 |
@@
| CAE [11] | arXiv 2022 | 300 | 1600 | 1600 | Pixel+Feat | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE [22] | arXiv 2022 | - | 1600 | 1600 | Pixel+Feat | - | 84.4 | - | 50.1 |
@@
| BootMAE [14] | ECCV 2022 | - | 800 | 800 | Pixel+Feat | - | 84.2 | 85.9 | 49.1 |
| CAN [31] | arXiv 2022 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
@@
| dBOT [29] | ICLR 2024 | - | 1600 | 1600 | Feature | - | 84.5 | 86.6 | 49.5 |
| CrossMAE [17] | arXiv 2024 | 800 | 800 | 800 | Pixel | 79.3 | 83.7 | 85.4 | |
@@ 5 Experiment
5.0.1 Pre-training setup.
We pre-train our framework and other SSL frameworks through DTM on ImageNet-1K. We train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base models [35] with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4 following [5, 34]. Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. For the hyperparameters of DTM, we used 2 for , 1 for , 147 for , and 14 for all . We also pre-train various SSL frameworks through our DTM with the same fundamental setups. More detailed setups for pre-training are detailed in Appendix. All models are pre-trained using 8 V100-32GB GPUs.
| Framework | Method | FT () |
| MAE + CLIP | 82.6 | |
| Feature MIM | MAE + CLIP + DTM | 83.2 (+0.6) |
| (CLIP teacher) | BEiT v2 | 84.2 |
| BEiT v2 + DTM | 84.4 (+0.2) | |
| Image-level SSL | BYOL | 81.7 |
| BYOL + DTM | 82.1 (+0.4) |
| Method | ViT-S | ViT-B | ViT-L |
| MIM (Baseline) | 82.7 | 84.3 | 86.4 |
| DTM | 83.2 | 85.4 | 86.7 |
| Gain | (+0.5) | (+1.1) | (+0.3) |
5.0.2 Fine-tuning setup.
We fine-tune our pre-trained models on ImageNet-1K [39] by default following the standard protocol [19, 34]. Pre-trained ViT-S, ViT-B, and ViT-L are fine-tuned for 300, 100, and 50 epochs, respectively. Optimization is performed with AdamW using a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. Additional fine-tuning details are listed in Appendix.
5.1 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K [39] datasets. The comparisons include supervised learning and SSL methods with various supervision signals. When the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for a fair comparison. Table 3 reports the fine-tuning accuracies of ViT-S/B/L backbones. We observe that our MIM pre-trained by DTM achieves 83.2%, 85.4%, and 86.7% top-1 accuracies with ViT-S/16, ViT-B/16, and ViT-L/16, respectively, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP [48], DeepMIM [38], and BEiT v2 [34] by 1.0p, 0.6p, and 0.4p on ViT-B/16, respectively. Moreover, our method outperforms other methods that leverage diverse supervision, demonstrating our method’s superiority among self-supervised learning methods.
5.2 Effectiveness of Our Method
5.2.1 Applicability on SSL frameworks.
We apply DTM on BEiT v2, MAE with the CLIP teacher [35], and BYOL [18] to verify its applicability to various SSL frameworks. As shown in Table 4, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability to various frameworks beyond our baseline.
5.2.2 Scalability.
We further study the scalability of DTM. We compare the fine-tuning accuracies of our method to its baseline, which is pre-trained without DTM. We pre-train ViT-S/16, ViT-B/16, and ViT-L/16 with a resolution of for each method and evaluate the performance gains in fine-tuning accuracies. As shown in Table 5, DTM improved 0.4p, 1.1p, and 0.3p for ViT-S/16, ViT-B/16, and ViT-L/16, respectively, verifying its general efficacy across the scales.
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
5.2.3 Further Applicability on SLIP.
5.2.4 Efficiency.
To verify the efficiency of our method, we compared Table 7 shows that bipartite matching is efficient and effective, significantly boosting the accuracy (+1.1%p) with only a 1% speed loss. While K-Means [30] also exhibits considerable improvements, it significantly degrades the training speed. Layer-wise K-Means shows accelerate the K-Means method by aggregating tokens within layers at the cost of degraded representations, leading to lower accuracy.
| Case | Throughput (image/s) | FT () |
| Baseline | 1458 | 84.3 |
| Layer-wise K-Means clustering | 1265 | 85.1 |
| DTM (K-Means clustering) | 489 | 85.4 |
| DTM (Bipartite matching) | 1446 | 85.4 |
5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performance on ADE20K [55] to verify the transferability of our pre-trained model. We follow the training and evaluation protocol [19]; the models are fine-tuned for 160K iterations using UperNet [51] with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 3 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method effectively enhances discriminability for dense prediction tasks.
| Method | iNat 2018 | iNat 2019 | iNat 2021-mini |
| Baseline | 75.0 (74.6±0.6) | 81.1 (79.8±1.0) | 75.8 (75.2±0.6) |
| DTM (Ours) | 78.5 (77.4±0.6) | 81.9 (81.2±0.6) | 78.4 (77.5±0.6) |
@@ 5.3 ADE20K Semantic Segmentation
5.4 Transfer learning
5.4.1 iNaturalist datasets.
5.4.2 Fine-Grained Visual Classification (FGVC) datasets.
| Case | Fine-tuning () |
| 84.5 | |
| 84.6 | |
| Smoothed | 84.4 |
| Cosine distance (Cos) | 84.8 |
| Case | Fine-tuning () |
| 84.7 | |
| 84.8 | |
| 84.7 |
| Method | Dynamic | FT () | |
@@ 5.4 Transfer learning
5.5 Ablation study
We conduct ablation studies on loss functions, the range of token morphing steps, and target normalization. We also study the impacts of the number of morphing schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K [39].
5.5.1 Loss function.
We compare various options for the loss function in our method. We compared , , smoothed , and cosine distance. As shown in Table LABEL:tab:ablation_loss_ftn, the model pre-trained using cosine distance outperforms the models with other distance functions.
5.5.2 Range of token morphing iterations.
We compare fine-tuning accuracies of the pre-trained models while varying the ranges that randomly samples the number of morphing iterations in Table LABEL:tab:ablation_step. While our DTM works for all the sampling ranges, K=14 works best.
@@ 5.6 Analysis
5.6.1 Ablation study on the importance of dynamics.
Table 11 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ ViT-B/16 [15] with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K [39]. For token morphing without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 11, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, demonstrating the importance of dynamic mechanism in DTM.
5.6.2 Faster convergence of DTM.
We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 5(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 5(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss than the baseline model, verifying better convergence. Finally, Fig. 5(c) verifies the consistent effect of DTM across various pre-training epochs.
@@ 6 Conclusion
We have introduced a novel masked image modeling method based on token morphing to address the negative impacts of spatially inconsistent target representations during pre-training. We have first analyzed the existence and impacts of spatial inconsistency in target representations. Specifically, we qualitatively observed spatial inconsistency among tokens from pre-trained models despite proximity and contextual similarity. We then investigated the impacts on downstream tasks, including zero-shot/linear classification. We have further validated that context-aware token aggregation methods enhance the pre-training capability of the target tokens, while arbitrary aggregation, like downsampling, tends to disrupt them. Based on the observations, we have proposed Dynamic Token Morphing (DTM), which dynamically aggregates contextually associated tokens with a random number of matching tokens. Subsequently, our method aligns the representations of morphed tokens with the representations of their corresponding morphed targets. Our extensive experiments verified its general applicability, scalability, and superiority. We further validated its transferability, and subsequent analyses supported our method. Our study showcased that a token aggregation method could boost representation learning performance, with the potential for broader future applications and research directions.
@@
Limitation. Despite the potential of our method, we have verified its applicability only up to ViT-L/16. Resource limitations prevented us from performing more experiments on larger-scale models such as ViT-G.
Appendix
@@
This appendix includes additional experimental analyses of our proposed method. First, we present additional examples of spatial inconsistency among patches from pre-trained models in §0.A. Then, we validate the applicability of superpixel algorithms [2] into DTM in §0.B. We also conduct ablation studies on number of morphing schedules and effects of randomness in the number of morphing tokens, randomness in gradual token morphing, and target normalization in §0.C. Finally, §0.D provides implementation details for both pre-training and fine-tuning on ImageNet-1K [39], as well as fine-tuning on ADE20K [55].
Appendix 0.A More Examples on Spatial Inconsistency
@@
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize token morphing results and correct maps of the patch-wise prediction results with and without token morphing using CLIP [35]. Specifically, the correct maps provide information about both correct and incorrect predictions. Patches with correct predictions are highlighted in yellow, while incorrectly predicted patches exhibit results with monotone colors. As shown in Fig. A, patch representations without token morphing reveal spatially inconsistent patch-wise predictions compared to the prediction results with token morphing, which reveals the spatially inconsistent representations among patches. In addition, predictions with token morphing exhibit significantly more correctly predicted patches than predictions without token morphing.
Appendix 0.B Compatibility with Superpixel Algorithms
We perform additional experiments to use the concept of superpixels [2, 8] in ours, both directly to tokens (layer-wise superpixel) and can be combined with DTM. The layer-wise superpixel method uses constant numbers of superpixels and iterations. Table A shows superpixel-based methods enjoy notable gains but do not exceed the bipartite matching one. While the layer-wise superpixel method also utilizes superpixels, the layer-wise token aggregation across the encoder layers bipartite matching approach risks harming intermediate representations during encoding
| Method | Fine-tuning () |
| Baseline | 79.5 |
| Layer-wise superpixel [2] | 86.7 |
| DTM (Superpixel clustering) [8] | 87.1 |
| DTM (Bipartite matching) | 87.9 |
| Method | Case | IN-1K Fine-tuning | IN-100 Fine-tuning |
| Baseline | 83.5 | 79.5 | |
| DTM | 1 | 84.8 | 87.6 |
| 2 | 84.9 | 87.9 | |
| 3 | 84.8 | 87.8 | |
| 4 | 84.8 | 87.8 |
| Method | Configurations | Fine-tuning () |
| Baseline | 79.5 | |
| DTM (ours) | Fixed number of morphing tokens | 87.4 |
| Single step morphing | 87.7 | |
| + Target normalization [19] | 87.7 | |
| Default | 87.9 |
| Hyperparameters | ViT-S/16 | ViT-B/16 | ViT-L/16 |